OpenStudy (mtalhahassan2):

what is the vertex of x square -3

OpenStudy (mtalhahassan2):

0,-3

OpenStudy (mtalhahassan2):

how can we graph this?

OpenStudy (aum):

Are you familiar with the graph of y = x^2 ?

OpenStudy (mtalhahassan2):

Is we have to used the table of value?

OpenStudy (mtalhahassan2):

yeah I am

OpenStudy (mtalhahassan2):

tell me the easier way

OpenStudy (aum):

If you are familiar with y = x^2, then the graph of y = x^2 - 3 is just the graph of y = x^2 shifted DOWN by 3 units.

OpenStudy (aum):



OpenStudy (aum):

If you are asked to do it by table then put x = -3, -2, -1, 0, 1, 2, 3 and calculate the y values. Plot the point and draw a graph: x f(x)=x^2 - 3 -3 (-3)^2-3 = 6 -2 -1 0 1 2 3

OpenStudy (mtalhahassan2):

like the value of original be 0,0, 1,1 2,2

OpenStudy (mtalhahassan2):

which way is more easier

OpenStudy (mtalhahassan2):

the table of value or just drawing the original graph then changing the points

OpenStudy (aum):

No. That will be for y = x^2 Here it is y = x^2 - 3 So you have to plug in each x value and calculate the y-value. Depends on what they are expecting. If they just want a general sketch the first method is easier. If they want an accurate plot the second method of table has to be used.